OpenStudy (anonymous):

I don't really get the options on this one I just got 6 (plus or minus) 0 when I worked it out... Which of the following equations is the result of completing the square on x^2 - 6x - 9 = 0? a) (x - 3)^2= 0 b) (x - 3^)2= 9 c) (x - 3)^2= 18

OpenStudy (johnweldon1993):

Well in order to complete the square...you would add 9 to both sides... x² -6x = 9 Now take your 'b' value (-6) divide it by 2...and square the result....you add this to both sides of the equation... so -6/2 = -3...and -3² = 9...so now we have x² - 6x + 9 = 9 + 9 x² - 6x + 9 = 18 And when you factor the left hand side you get?

OpenStudy (johnweldon1993):

So general rules for completing the square... 1) Get into the form * Ax² + Bx = C * 2) Take 'B' and divide it by 2...and square the result... 3) Add that result to both sides of the equation 4) Factor the left hand side and *although you dont need to here* 5) Take the square root of both sides 6) Solve for 'x'

OpenStudy (anonymous):

well to answer the other question 9? but my diddy told me I was wrong

OpenStudy (anonymous):

(x - 3)^2= 18

OpenStudy (johnweldon1993):

Right ...that is correct...(x - 3)² = 18 is what I get too

OpenStudy (anonymous):

yippee! thanks! @johnweldon1993

OpenStudy (johnweldon1993):

No problem! :)